solve the following equation for x and y :    log of mod of x+y with base 100 =1/2 and log of y with base 10 -log of mod x with base 10 =log of4 with base 100
 

Grade:11

1 Answers

noogler
8 years ago
x=10/3,y=20/3.........or.............x=-10/3,y=-20/3